John runs a computer software store. Yesterday he counted 140 people who walked by the store, 63 of whom came into the store. Of the 63, only 25 bought something in the store.
    (a) Estimate the probability that a person who walks by the store will enter the store. (Round your answer to two decimal places.)
1

    (b) Estimate the probability that a person who walks into the store will buy something. (Round your answer to two decimal places.)
2

    (c) Estimate the probability that a person who walks by the store will come in and buy something. (Round your answer to two decimal places.)
3

    (d) Estimate the probability that a person who comes into the store will buy nothing. (Round your answer to two decimal places.)
4
Answer #1

a) 0.45

b) 0.39

c) 0.17

d) 0.60
Answer #2

(a) probability that a person who walks by the store will enter the store =

(No. of persons who enteres the store)/(No. of persons who walks by the store)

= 63/ 140

= 0.45


(b) probability that a person who walks into the store will buy something

= (No. of perosns who bought something) / (No. of persons who walks into the store)

= 25/63

= 0.396825..........

= 0.40


(c) probability that a person who walks by the store will come in and buy something

= (No .of persons who bought something) / (No. of persons who walks by)

= 25 / 140

= 0.1785.....

= 0.18


(d) probability that a person who comes into the store will buy nothing.

= (No. of persons who came in but bought nothing) / (No. of persons who cam into the store)

= (63-25)/63

= 0.6031...

= 0.60